Problem: x^2+6x+5=0
1)Factor
2)Completing Square

3) Quadratic Formula

Answer by stanbon(75887)

Problem: x^2+6x+5=0
1)Factor
(x+5)(x+1) = 0
x = -5 or x = -1
=======================
2)Completing the Square
x^2 + 6x + 9 = -5+9
(x+3)^2 = 4
---
x+3 = 2 or x+3 = -2
x = -1 or x = -5
=======================================
3) Quadratic Formula
x = [-6 +- sqrt(36-4*5)]/2
---
x = [-6 +- sqrt(16)]/2
---
x = [-6 +- 4]/2
---
x = (-6-4)/2 = -5
OR
x = (-6+4)/2 = -1
=====================
Cheers,
Stan H.
=====================